OpenStudy (anonymous):

Find the slope of the tangent line to the graph of f at the given point. f(x) = square root of x, at (36, 6) A. 1/3 B. 12 C. 1/12

OpenStudy (anonymous):

derivative of \(f(x)=\sqrt{x}\) is \(f'(x)=\frac{1}{2\sqrt{x}}\) memorize that one

OpenStudy (anonymous):

\[f(x)=\sqrt{x}\] \[f'(x) = \frac{1}{2\sqrt{x}}\] \[evaluate \frac{1}{2*\sqrt{36}} = \frac{1}[\]

OpenStudy (anonymous):

math teachers love to put it on quizzes and test, so while your classmates waste their time using the power rule, you are done in a second

OpenStudy (anonymous):

1/12

OpenStudy (anonymous):

Thank you everyone

OpenStudy (anonymous):

what @Fonzy said and don't forget "the derivative of root x is one over two root x" all day every day

OpenStudy (anonymous):

Here's a more complete step so you can see if you are a visual type, \[f(x) =\sqrt{x} \] \[f(x) = x^{\frac{1}{2}} \] now apply the power rule.\[nx^{n-1}\] \[f'(x) = \frac{1}{2}x^\frac{-1}{2}\] \[f'(x) = \frac{1}{2\sqrt{x}}\]